# LoRAMoE: Revolutionizing Mixture of Experts for Maintaining World Knowledge in Language Model Alignment

### 18 Dec 2023 v2

@@ Abstract
Supervised fine-tuning (SFT) is a crucial step for large language models (LLMs), enabling them to align with human instructions and enhance their capabilities in downstream tasks. When the models are required to align with a broader range of downstream tasks, or there is a desire to notably improve the performance on a specific task, a substantial increase in fine-tuning data often emerges as the solution. However, we find that large-scale increases in instruction data can disrupt the world knowledge previously stored in the LLMs, i.e., world knowledge forgetting. In this paper, we introduce LoRAMoE to address the above challenge. The LoRAMoE is a plugin version of Mixture of Experts (MoE). The plugin form ensures the integrity of world knowledge by freezing the backbone model during the training phase. We then propose the use of localized balancing constraints to coordinate parts of experts for task utilization, meanwhile enabling other experts to fully leverage the world knowledge stored in the models. Experimental results demonstrate that LoRAMoE can reasonably coordinate experts based on data type during inference, and even dramatically increasing instruction data does not result in knowledge forgetting. Moreover, LoRAMoE provides additional benefits for the performance of downstream tasks, indicating the potential of our approach for multi-task learning.
@@ 1 Introduction
Large Language Models (LLMs) (Touvron et al., 2023; Muennighoff et al., 2022) have demonstrated remarkable capabilities in a variety of tasks. Supervised fine-tuning (SFT) of the models to align them with human instructions is a crucial step in unleashing their full potential (Chung et al., 2022; Ouyang et al., 2022). Although some works (Zhou et al., 2023; Cao et al., 2023) have indicated that models can follow human instruction well with a little fine-tuning data, increasing the amount of data is a straightforward solution when the variety of tasks expands, or when enhanced performance on a specific task is required (as shown in the left of Figure 1).
However, the large-scale increase in fine-tuning data brings new challenges. Specifically, we observe a notable decline in performance on the Closed-Book Question Answering (CBQA) dataset (e.g., TriviaQA (Han et al., 2019), Natural Questions (Kwiatkowski et al., 2019)) when there is a substantial increase in the amount of fine-tuning data, as shown in the blue group of lines on the right of Figure 1. We hypothesize that this significant decline in performance may be related to the collapse of world knowledge (Touvron et al., 2023) (i.e., parametric knowledge Neeman et al. (2022)) previously learned and stored in the pre-trained models (Petroni et al., 2019; Yu et al., 2023). This hypothesis is demonstrated in two steps. First, we verify that the CBQA dataset relies on the world knowledge stored in the models for making inferences. Second, we demonstrate that the substantial decrease in performance on the CBQA dataset is attributed to the fact that large-scale fine-tuning can markedly change the model’s parameters (as shown in the right of Figure 1), leading to the destruction of world knowledge (i.e., knowledge forgetting). Overall, in vanilla supervised fine-tuning, there is a contradiction between simultaneously improving performance on downstream tasks and maintaining world knowledge of LLMs.
Findings: Vanilla SFT with massive data can lead to forgetting parametric knowledge stored in LLMs.
Figure 1: (Left) When the number of fine-tuning data increases from 100,000 to 3 million, the performance of many tasks is significantly improved. (Right) With the amount of instruction data increasing, fine-tuning (training for an epoch using the same set of hyperparameters) can continually change model parameters (shown as the red line), resulting in a decline in performance on the benchmarks that measure world knowledge. The details of training implementation can be seen in Section 2.1.
An ideal solution is to delineate a specific region within the model dedicated to storing world knowledge, akin to how the hippocampus in the human brain is specialized for memory (Treves & Rolls, 1994; Voss et al., 2017). Such a structure allows the parameters in this part to be frozen and thus protected during the fine-tuning process, thereby keeping world knowledge from being disrupted. However, due to the black-box nature of large language models (Sun et al., 2022; Kaddour et al., 2023), identifying regions of the world knowledge of LLMs is highly challenging (Feng et al., 2023). Instead of identifying these regions, another solution is to keep all the parameters of the model by plugin-based fine-tuning. Specifically, it fine-tunes the model by freezing all the parameters of the backbone model and binding parameter changes to an additional network. With a complete backup of the backbone model’s parameters available, the world knowledge stored in it is technically recoverable.
However, fine-tuning with a single plugin is similar in form to direct fine-tuning (Ding et al., 2022; He et al., 2021b). Consequently, the issue of knowledge forgetting persists. Mixture of Experts (MoE) (Jacobs et al., 1991) is an architecture that introduces multiple experts, where data with different characteristics are routed to the corresponding experts for customized processing (Du et al., 2022; Shazeer et al., 2016). Drawing on this idea, we hope to introduce multiple plugins as experts, allowing a part of them the opportunity to access the backup, while another part can perform downstream tasks.
In this paper, we propose LoRAMoE, to alleviate world knowledge forgetting and simultaneously enhance the LLMs’ capabilities of solving downstream tasks. LoRAMoE is a plugin version of MoE. It changes the model’s architecture by adding multiple parallel plugins as experts (i.e. LoRA Hu et al. (2021)) in each feed-forward layer and connecting them with routers. We then propose the use of localized balancing constraint to split the experts in each LoRAMoE layer into distinct groups. Specifically, one group is dedicated to downstream tasks, while the other focuses on aligning world knowledge within the backbone model with human instructions to alleviate knowledge forgetting. Additionally, localized balancing constraint also balances the importance of all experts within the same expert group, which prevents only a few experts in the same group from being valued by the routers. It enables several experts to collaborate, improving capabilities of solving downstream tasks.
Experiment results show that LoRAMoE can effectively keep the world knowledge in language models from being disrupted by the large-scale fine-tuning. Further, we confirm the effectiveness of LoRAMoE on capability localization at an interpretable level by visualizing the expert weight for tasks. Our observations reveal when finishing world knowledge benchmarks, the router pays more attention to the output of experts specifically handling these tasks. Conversely, for other downstream tasks, the router focuses on experts from another group. LoRAMoE effectively resolves the conflict by fostering collaboration among experts. Besides, the experiment result shows that learning on various downstream tasks also benefits from our method, implying the prospect of our method in multi-task learning.
Our contributions can be summarized as follows:
@@
    We find that significantly increasing the amount of supervised fine-tuning data can severely impair the world knowledge inside the LLMs, due to its great modification to their parameters. This indicates that maintaining the world knowledge inside the LLMs is in conflict with the large-scale addition of downstream fine-tuning data.
@@
    We introduce a new trainable plugin for LLMs, LoRAMoE, which is similar in architecture to the Mixture of Experts. LoRAMoE can automatically route different types of data to respective experts during the SFT phase without interfacing with the original parameter of LLMs. LoRAMoE employs localized balancing constraint in training, enhancing expert group specialization and internal balance. It partitions experts into two groups to learn tasks and align world knowledge with human instructions, reducing knowledge forgetting.
@@
    We demonstrate the effectiveness of LoRAMoE through extensive experiments. We can maintain a stable knowledge in the model when scaling up fine-tuning data, during which performance in other tasks has also seen considerable improvement. Our method is further evidenced by the visualization of expert utilization.
2 Conflict between Expanding fine-tuning data and Retention of World Knowledge in LLMs
In this section, we conduct SFT tasks on the LLM with vast and varied datasets. We find that the world knowledge inside the LLM is severely compromised during the expansion of SFT.
2.1 Implementation
Datasets. To understand the impact of large-scale SFT on the world knowledge of the LLM, we construct a large-scale dataset that includes a variety of tasks. Namely, there are seven tasks, Closed-Book Question Answering (CBQA), coreference resolution, Natural Language Inference (NLI), abstract summarization, multi-lingual translation, reading comprehension, and text classification. We augmented the training datasets to 5 million through data augmentation methods. More details about our composition of fine-tuning data can be seen in the Appendix A.1.
Base Model. We utilize LLaMA-2-7B (Touvron et al., 2023) as the base model, considering it stands out as one of the most notable and widely used open-source LLMs in current academia.
Evaluation. For evaluation of the world knowledge, we use CBQA as a key benchmark of the model’s world knowledge, with reference to the previous work (Touvron et al., 2023; Petroni et al., 2019). Notably, considering previous work that has noted train-test overlap in CBQA datasets (Lewis et al., 2020), we elaborately select parts of the CBQA dataset without train-test overlap for our testing set, namely Filtered NQ and Filtered TriviaQA, to analyze the world knowledge of models better. For evaluating the performance on other downstream tasks, we utilize the opencompass [^1] https://opencompass.org.cn/ framework to run the evaluation process on the aforementioned tasks.
2.2 The Expansion of fine-tuning data Leads to the Knowledge Forgetting inside the LLMs
@@
During the expansion of fine-tuning data, we observed a diverging trend in the performance across two types of tasks, as can be seen in Figure 2:
On some tasks, such as summarization, NLI, machine translation, etc., the performance fine-tuned model initially increased magnificently and stabilized at a promising level. However, there was a catastrophic decline in the model’s performance on the benchmark measuring knowledge capability(e.g. TriviaQA, NQ, Hotpot QA), even much lower than the baseline. Notably, with the training data expanding, a contiguous decline can be witnessed. Besides, in the filtered test set, the collapse happens earlier on the filtered test set than the original one.
Further, we dissect the reasons behind the decline of the performance in world knowledge benchmarks with the expansion of fine-tuning data.
The performance on world knowledge benchmarks highly relies on the knowledge and skills learned during pre-training phase.
To investigate the relationship between world knowledge benchmarks and the knowledge embedded in pre-trained models, we conduct fine-tuning solely on the CBQA dataset with 250k samples and run evaluation on the test sets without train-test overlap.
The results in Figure 3 shows that: the performance on these benchmarks can be significantly enhanced through naive training, however, the first one percent of the training process (approximately 1000 samples) contributes to most of the boost, and further increases in the training sample do not actually improve the performance to any great extent.
Actually, this phenomenon is reasonable. In the early stages of fine-tuning, the model quickly learns to follow human instructions through training data and align the world knowledge already stored in the model with the instructions (Zhou et al., 2023), thereby increasing the performance of CBQA. However, due to the limited overlap between the training and testing data, the knowledge in the test set is difficult to incorporate during training, so more samples do not improve performance.
Therefore, the model’s capability for completing the world knowledge benchmark is highly dependent on the knowledge and skills acquired during pre-training phase. Given this, it is naturally assumed that the markedly diminished performance of the model on CBQA stems from the disruption of knowledge stored in the LLM due to large-scale instruction tuning.
SFT process with large-scale instruction data disrupts the stored knowledge in LLMs, thus leading to knowledge forgetting.
To verify this hypothesis, we employed two distinct datasets for sequential fine-tuning of the model. Initially, the model was fine-tuned using instruction data excluding the CBQA segment. Subsequently, we further fine-tuned the model with the CBQA dataset that had been segregated before.
The experimental results are presented in Table 1, where we can observe a notable degradation in the knowledge capabilities of the fine-tuned model and its performance was inferior compared to the original LLM. This indicates that the world knowledge within the model was compromised during the first stage of fine-tuning, resulting in the model’s inability to forge the alignment between human instructions and the already disrupted knowledge in the subsequent stage of fine-tuning.
Tasks Baseline
@@
Table 1: From left to right are the performance of LlaMA-2-7B, the model fine-tuned solely on CBQA datasets and the model first fine-tuned on 3 million instruction datasets excluding CBQA then continue-trained on CBQA datasets. After continuing to fine-tune on CBQA dataset, the model after a large-scale SFT still fails to improve its knowledge-answering ability and falls far below the baseline.
Further, we find that there is a massive change in the LLM’s parameter during the expansion of fine-tuning data, as can be seen in the right part of Figure 1. As the previous research has documented that models store knowledge within their parameters during the pre-training process (Petroni et al., 2019; Roberts et al., 2020; AlKhamissi et al., 2022), this further indicates the destruction of the knowledge stored in the parameters during large-scale fine-tuning, which results in the knowledge forgetting.
Overall, enhancing the instruction following capability and facilitating the performance on various downstream tasks of LLM through large-scale vanilla SFT inherently conflicts at the parameter level with retaining the world knowledge stored in the model.
3 Methodology
Enhancing the language model’s capabilities in various tasks while retaining its world knowledge is crucial for applying them across various downstream tasks. However, as previously mentioned, these two objectives conflict at the parameter level with traditional methods. To achieve this goal, we propose LoRAMoE, an LLM adapter employing the Mixture of Experts (MoE) approach, designed to partition different capabilities within distinct sections of LoRA (Low-Rank Adaptation). In this section, we initially provide a concise overview of the traditional MoE and LoRA methodologies. Subsequently, we delve into how LoRAMoE ingeniously amalgamates the methodologies of MoE and LoRA. This integration effectively leverages the strengths of both methods and capably addresses the conflict issue outlined in section 2.
3.1 Preliminaries
3.1.1 Mixture of Experts
The Mixture of Experts significantly scales up model parameters without correspondingly increasing computational efforts. For transformers-based large language models, MoE supplants the conventional feed-forward neural network layer in each transformer block with an MoE layer (Shazeer et al., 2016; Fedus et al., 2021; Lepikhin et al., 2020). This MoE layer is composed of $N$ parametrically identical and independent feed-forward neural networks ${\{E_{i}\}}_{i=1}^{N}$ as the experts, coupled with a gating function $G(\cdot)$ as the router. The router is used to model the probability distribution that governs the weights of outputs from these expert networks. Formally, for the output $h$ of the attention layer in any given block, the output $y$ of the MoE layer can be mathematically represented as follows:
$y=\sum_{i=1}^{N}{G(h)_{i}E_{i}(h)}$ (1)
where $E_{i}(h)$ and $G(h)_{i}$ denote the output and the corresponding weight of $i$-th expert in the MoE layer, respectively. The router $G(\cdot)$ can be written as follows:
$G(\cdot)=Softmax(hW_{g})$ (2)
where $W_{g}$ is the trainable weight matrix for router $G(\cdot)$.
The language model based on MoE architecture facilitates the acquisition of varied competencies by its experts and their focus on distinct capabilities, achieved through the mechanism of routing (Du et al., 2022; Riquelme et al., 2021; Bao et al., 2022). Our work draws upon this MoE paradigm and we fine-tuned the experts to strategically distribute competencies, aiming to address the knowledge-forgetting issue. To clarify, we freeze the parameters of the backbones and substitute the feed-forward neural network layer of experts with LoRA to enhance the efficiency of the training process.
3.1.2 Low-Rank Adaption
LoRA (Low-Rank Adapter) has been demonstrated to be an effective and efficient way to adapt Pre-trained Models to specific tasks (Hu et al., 2021). Formally, for a pre-trained matrix $W$ (e.g., attention-k) with the $W_{0}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$, LoRA update the $W$ with a low-rank decomposition:
$W+\Delta W=W+BA,A\in\mathbb{R}^{d_{\text{in}}\times r},B\in\mathbb{R}^{r\times d_{\text{out}}}$ (3)
where r represents LoRA rank. The forward process with LoRA is as follows, during which the $W$ is frozen:
$h=Wx+\frac{\alpha}{r}BAx$ (4)
where $\alpha$ represents a scaling factor that adjusts the magnitude of the changes on the original $W$ made by LoRA modules. It is worth noting that, the forward process of LoRA is different from LoRAMoE. In the latter, LoRA is used as the expert network within the MoE architecture, primarily to facilitate accelerated training and optimization.
3.2 LoRAMoE
@@
3.2.1 Architecture
Our goal is to tackle the challenge of conflict in the expansion of instruction data with the maintenance of world knowledge inside LLM during the fine-tuning phase. The left of Figure 4 illustrates the forward process of the standard MoE architecture. In the MoE-based architecture, the router function focuses on different experts according to the instruction data categories, allowing them to divide their labor to complete the forward process (Jacobs et al., 1991).
Taking inspiration from that discovery, as shown on the left side of Figure 4, we attempt to extend the pre-trained language model architecture into an MoE-like architecture to tackle the aforementioned conflict, at the same time retaining its original strong capabilities of language model. Specifically, we freeze the parameters of the backbone model, allowing experts within LoRAMoE the opportunity to leverage the existing world knowledge in the base model. For the feed-forward neural network layer in each transformers block, we define several parallel trainable experts connected by the router. Simultaneously we replace the fully-connected layer of the expert with a low-rank form to improve training and inference efficiency.
Formally, for the traditional transformers architecture, the forward propagation process of each decoder block can be simplified as follows:
$f(x)=x+f_{\text{FNN}}(x)$ (5)
where $f_{\text{FFN}}(\cdot)$ stands for the feed-forward neural network block and $x$ denotes the input of the FFN block. The matrix operation of the linear layer in the FNN block can be expressed as:
@@ 3.2.1 Architecture
where $W_{0}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ represents the parameter matrix and $\Delta W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ denotes the parameter update in the training phase. $o$ is the output of the linear layer with dimension $d_{\text{out}}$.
To address the conflict issue delineated in Section 2 during the fine-tuning phase, we adopted a strategy distinct from optimizing a single matrix with the entire instruction dataset. Specifically, we substituted the linear layer with the MoE architecture. This modification permits the experts within the MoE layer to collaborate and learn the updated matrix $\Delta W$. Consider the MoE layer comprising $N$ experts and the set of experts denoted as $\{E_{i}\}_{i=1}^{N}$, the forward process of the MoE layer can be mathematically expressed as follows:
@@
where $E_{i}(\cdot)$ and $G(\cdot)=Softmax(xW_{g})$ represent the $i$-th expert and the router in the MoE layer, respectively. The $W_{g}$ is the trainable parameter matrix of route function. Through this approach, the experts and routing function work in tandem during the training phase, enabling the experts to develop varied capabilities and efficiently handle diverse types of tasks.
Concurrently, Low-Rank Adaptation has been proven to be both effective and efficient in the fine-tuning of pre-trained language models (Wang et al., 2023; Liu et al., 2022; Pan et al., 2022). To enhance the efficiency and resource conservation of the fine-tuning process, we adapt the parameter matrix of the experts in the MoE layer into a low-rank format. Specifically, the matrix $\Delta W_{E}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ of single expert $E(\cdot)$ in the LoRAMoE layer can be written as follows:
@@
where $A\in\mathbb{R}^{d_{\text{in}}\times r}$, $B\in\mathbb{R}^{r\times d_{\text{out}}}$, and the rank $r\ll\min(d_{\text{in}},d_{\text{out}})$. LoRA contributes to a significant reduction in the parameters of the experts network, thereby enhancing efficiency and saving costs during the fine-tuning process. Overall, the forward process of the modified linear layer (i.e., LoRAMoE Layer) in the feed-forward neural network layer can be represented as:
$o=W_{0}x+\gamma\sum_{i=1}^{N}G(x)_{i}E_{i}(x)=W_{0}x+\frac{\alpha}{r}\sum_{i=1}^{N}\omega_{i}\cdot B_{i}A_{i}x$ (9)
@@
where $\omega_{i}$ denotes the attention weight of $i$-th expert in the LoRAMoE layer and $\alpha$ is the constant hyper-parameter, approximately equivalent to the learning rate (Hu et al., 2021).
@@
Figure 5: The coefficient of variation for the experts of the unconstrained LoRAMoE progressively escalates and sustains at a high level (i.e., approximately three, similar to the phenomenon observed at Shazeer et al. (2016)), signifying the prolonged predominance of a limited number of experts.
Figure 6: The experts are segregated into two categories: one of which concentrates on learning massive tasks, while another focuses on aligning the world knowledge inside the LLM and human instruction. The routing mechanism assigns similar importance to the experts within a group, and it selectively activates the group more specialized for the given data type.
3.2.2 Mixed Distribution Dilemmas for Expert Balancing
$L(\mathbf{X})=$ $\prod_{x\in\mathbf{X}_{1}}\left(m\times\mathcal{N}\left(x;\mu^{{}^{\prime}}_{1},\sigma^{{}^{\prime}2}_{1}\right)+(1-m)\times\mathcal{N}\left(x;\mu^{{}^{\prime}}_{2},\sigma^{{}^{\prime}2}_{2}\right)\right)\times$
$\prod_{x\in\mathbf{X}_{2}}\left(m\times\mathcal{N}\left(x;\mu^{{}^{\prime}}_{1},\sigma^{{}^{\prime}2}_{1}\right)+(1-m)\times\mathcal{N}\left(x;\mu^{{}^{\prime}}_{2},\sigma^{{}^{\prime}2}_{2}\right)\right),$ (10)
$\frac{\partial\log L(\mathbf{X})}{\partial\mu_{1}^{\prime}}$ $=\sum_{x\in\mathbf{X}_{1}\cup\mathbf{X}_{2}}\frac{\partial}{\partial\mu_{1}^{\prime}}\log\left(m\times\mathcal{N}_{1}(x)+(1-m)\times\mathcal{N}_{2}(x)\right)$
$=\sum_{x\in\mathbf{X}_{1}\cup\mathbf{X}_{2}}\left(\frac{\left(x-\mu_{1}^{\prime}\right)}{\sigma_{1}^{\prime 2}}\right)\times\frac{m\times N_{1}(x)}{m\times N_{1}(x)+(1-m)\times N_{2}(x)},$ (11)
$\frac{d\log L}{dm}$ $=\frac{\partial\log L}{\partial\mu_{1}^{\prime}}\times\frac{d\mu_{1}^{\prime}}{dm}$
$=\left(\sum_{x\in\mathbf{X}_{1}\cup\mathbf{X}_{2}}\left(\frac{\left(x-\mu_{1}^{\prime}\right)}{\sigma_{1}^{\prime 2}}\right)\times\frac{m\times N_{1}(x)}{m\times N_{1}(x)+(1-m)\times N_{2}(x)}\right)\times\frac{p_{1}(\mu_{2}-\mu_{1})}{m^{2}}$
$\leq 0$ (12)
The inverse result can be derived similarly. Therefore, the best training error is achieved only when the mixing coefficient $m$ of the prior distribution is consistent with the actual sampling distribution weight $p_{1}$. In the next section, we will introduce a localized balancing constraint algorithm, which uses adaptive balancing coefficients to explore the optimal data distribution mixing coefficient.
3.2.3 Localized Balancing Constraint
Considering the aforementioned dilemmas, we introduce a strategic approach as depicted in Figure 6. Unlike previous efforts that solely focused on distributing the learning data among all experts evenly, LoRAMoE can methodically segregate the experts into two distinct groups, one of which concentrates on learning massive tasks, while another focuses on aligning the world knowledge with instructions (i.e. invoking the world knowledge in the base model to take respond to the human instructions). While all experts within the same group are balanced, this balance is not uniformly applied across groups.

### 05 Mar 2024 v3

# LoRAMoE: Alleviate World Knowledge Forgetting in Large Language Models via MoE-Style Plugin

@@ Abstract
Supervised fine-tuning (SFT) is a crucial step for large language models (LLMs), enabling them to align with human instructions and enhance their capabilities in downstream tasks. Increasing substantially instruction data is a direct solution to align the model with a broader range of downstream tasks or notably improve its performance on a specific task. However, we find that large-scale increases in instruction data can damage the world knowledge previously stored in LLMs. To address this challenge, we propose LoRAMoE, a novelty framework that introduces several low-rank adapters (LoRA) and integrates them by using a router network, like a plugin version of Mixture of Experts (MoE). It freezes the backbone model and forces a portion of LoRAs to focus on leveraging world knowledge to solve downstream tasks, to alleviate world knowledge forgetting. Experimental results show that, as the instruction data increases, LoRAMoE can significantly improve the ability to process downstream tasks, while maintaining the world knowledge stored in the LLM 111 The code will be made available upon publication..
LoRAMoE: Alleviate World Knowledge Forgetting in Large
Language Models via MoE-Style Plugin
[…]
Figure 1: (Left) With the number of fine-tuning data increases from 10K to 3M, the performance of many downstream tasks is significantly improved. (Right) With the amount of instruction data increasing, fine-tuning the language models results in a decline in performance on the benchmarks that measure their world knowledge, such as TriviaQA (Han et al., 2019), Natural Questions (Kwiatkowski et al., 2019). The details of training implementation can be seen in Section 2.1.
@@ 1 Introduction
Supervised fine-tuning (SFT) provides a pivotal technique to make large language models (LLMs) follow human instructions and improve their performance of downstream tasks (Chung et al., 2022; Ouyang et al., 2022). Although some studies (Zhou et al., 2023; Cao et al., 2023) indicate that LLMs trained on a little data can follow instructions well, increasing the amount of data is a straightforward way to enhance their ability to multiple downstream tasks or improve their performance on a specific task, as shown in the left of Figure 1.
However, the large-scale increase in instruction data can destroy the world knowledge stored in LLMs, as illustrated in the right of Figure 1. Specifically, as the amount of instruction data increases, we observe a notable decline in performance on Closed-Book Question Answering (CBQA) datasets, which are used to measure world knowledge in LLMs Touvron et al. (2023); Neeman et al. (2022). In the paradigm of supervised fine-tuning, the conflict between maintaining world knowledge inside LLMs and improving their performance on downstream tasks by scaling up instruction data has not been thoroughly examined.
In this paper, we propose LoRAMoE, a novelty framework for SFT, to enhance the models’ capability of solving downstream tasks, while alleviating world knowledge forgetting during the training phase. LoRAMoE is a Mixture-of-Experts-style (MoE-style) plugin, which introduces several low-rank adapters (LoRA Hu et al. (2021)) as experts and integrates them by using a router network. The router network automatically assigns weights to experts, which can improve the LLM’s performance on multiple downstream tasks.
To demonstrate the efficacy of our proposed method, we conduct extensive experiments across a range of downstream tasks. Experiment results show that LoRAMoE can significantly improve LLM’s ability to address the various downstream tasks by fine-tuning the model on a large amount of instruction data, while maintaining the world knowledge stored in the model. In addition, we further evaluate our method by visualizing the expert weight for tasks. The result indicates that LoRAMoE adequately alleviates world knowledge forgetting and achieves an improvement of models by fostering collaboration among experts. The main contributions of our paper are as follows:
@@
    We find that significantly increasing the amount of instruct data during the SFT phase can damage the world knowledge inside the LLMs. The need for improvement in downstream tasks by scaling up instruction data conflicts with maintaining the world knowledge inside the model.
@@
    We introduce LoRAMoE, a novelty framework for SFT, which introduces LoRAs as experts and integrates them by the router. LoRAMoE can enhance the model’s ability to address downstream tasks, while alleviating the world knowledge forgetting.
@@
    Extensive Experiments demonstrate the efficacy of our proposed approach in multi-tasks and mitigating the forgetting of world knowledge inside the model. The visualizing experiment shows that LoRAMoE can achieve an improvement by fostering collaboration among experts.
2 Motivation
In this section, we verify that a large-scale SFT can cause irreversible damage to world knowledge within the LLMs while improving the LLMs’ performance in various downstream tasks.
2.1 A Diverging Trend
We constructed a dataset containing seven categories of tasks with a total of five million training samples, and used it to conduct SFT on a Llama-2-7B model. The implementation details are described in Appendix A. During the expansion of fine-tuning data, we observed a diverging trend in the performance across two types of tasks, as shown in Figure 2:
Across downstream tasks such as summarization, Natural Language Inference (NLI), machine translation, and others, the performance of the fine-tuned model initially showed a magnificent increase and eventually stabilized at a promising level. However, when it comes to closed-book QA (CBQA) tasks that are used as world knowledge benchmark (Touvron et al., 2023; Neeman et al., 2022), the model’s performance catastrophically declines under the baseline Notably, with the training data expanding, a contiguous decline can be witnessed. Moreover, this decline will occur earlier if the test set is filtrated. Appendix B case with a larger dataset including more tasks shows an even steeper drop on world knowledge benchmarks, although performance remains competitive on others.
2.2 The Irreversible Knowledge Forgetting
In this section, we dissect the reason behind the decline on these world knowledge benchmarks during the expansion of fine-tuning data. We find this results from the occurrence of irreversible knowledge forgetting inside the LLM.
@@ 2.2 The Irreversible Knowledge Forgetting
The performance on world knowledge benchmarks highly relies on the knowledge and skills learned during pre-training phase. To investigate the relationship between the performance on world knowledge benchmarks and the knowledge embedded in pre-trained models (Petroni et al., 2019; Roberts et al., 2020; AlKhamissi et al., 2022), we conduct fine-tuning solely on the CBQA dataset with 250k samples and run evaluation on the test sets without train-test overlap. Results in Figure 3 show initial training boosts performance significantly, especially the first 1% (approximately 1k samples), with limited gains thereafter. This is because early fine-tuning aligns existing knowledge with new instructions, improving CBQA results. However, due to minimal training-testing data overlap, adding more samples doesn’t further enhance performance. Thus, a model’s benchmark success relies on world knowledge acquired from the pre-training.
Given this, it is naturally assumed that the diminished performance on knowledge benchmark stems from the damage of knowledge stored in the LLM due to large-scale instruction tuning. To verify the hypothesis, we sequentially fine-tuned a model using two datasets, first excluding CBQA data, then with CBQA data. Results presented in Table 1 show a great decline in knowledge capabilities versus the original LLM. This indicates that the world knowledge within the model was compromised during the first stage of large-scale fine-tuning, resulting in the model’s inability to forge the alignment between human instructions and the already destroyed knowledge in the subsequent stage of fine-tuning solely with CBQA.
To sum up, the pursuit of enhancing performance on downstream tasks through the expansion of training data conflicts the preservation of world knowledge within the model in vanilla SFT.
Task Name Baseline
@@
Table 1: Performance from left to right: LlaMA-2-7B, model tuned on CBQA, and model tuned on 3M instructions then on CBQA. Despite further tuning on CBQA, the large-scale SFT model’s knowledge-answering doesn’t improve, staying below the baseline.
@@
3 LoRAMoE
In this section, we elaborate on the methodological details of LoRAMoE, which is an MoE-style plugin and introduced Localized Balancing Constraint during the training phase to alleviate the world knowledge, as shown in Figure 4.
3.1 Architecture
The left of Figure 4 illustrates the forward process of the standard MoE architecture (Shazeer et al., 2016; Fedus et al., 2021; Lepikhin et al., 2020). In the MoE, the router assigns weights of experts according to the data, allowing them to divide their labor to complete the forward process (Jacobs et al., 1991). The key sight of LoRAMoE is that we freeze the backbone model to maintain world knowledge and introduce experts to leverage this knowledge to address tasks, while improving the performance on multiple downstream tasks. Additionally, we utilize the LoRA Hu et al. (2021) as the architecture of the expert to improve training and inference efficiency.
Formally, for the traditional transformers architecture, the forward propagation process of the feed-forward neural (FFN) network block can be simplified as follows:
$f(x)=x+f_{\text{FNN}}(x).$ (1)
The matrix operation of the linear layer in this forward propagation can be expressed as:
@@ 3.1 Architecture
where $W_{0}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ represents the parameter matrix of the backbone model and $\Delta W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ denotes the updated parameter during the training phase. For LoRAMoE, we replace the linear layer in the FFN block with the MoE-style plugin, which makes experts collaborate to address tasks. During the training phase, we freeze the backbone to maintain the world knowledge and only update $\Delta W$. Consider the LoRAMoE layer containing $N$ experts, which is denoted as $\{E_{i}\}_{i=1}^{N}$, the forward process of the layer can be mathematically expressed as follows:
@@
where $E_{i}(\cdot)$ and $G(\cdot)=Softmax(xW_{g})$ represent the $i$-th expert and the router in the LoRAMoE layer, respectively. The $W_{g}$ is the trainable parameter matrix of the route network. By this, the experts and the outer work in tandem, enabling the experts to develop varied capabilities and efficiently handle diverse types of tasks.
In addition, LoRA has been proven to be both effective and efficient for the SFT phase of LLMs (Wang et al., 2023a; Liu et al., 2022; Pan et al., 2022). To enhance the efficiency and resource conservation of the fine-tuning process, we replace the parameter matrix of the experts with a low-rank format. Specifically, the matrix $\Delta W_{E}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ of the expert $E(\cdot)$ in the LoRAMoE layer can be written as follows:
@@
where $A\in\mathbb{R}^{d_{\text{in}}\times r}$, $B\in\mathbb{R}^{r\times d_{\text{out}}}$, and the rank $r\ll\min(d_{\text{in}},d_{\text{out}})$. LoRA contributes to a significant reduction in the trainable parameters, thereby enhancing efficiency and saving costs during the fine-tuning process.
Overall, the forward process of the LoRAMoE layer replaced the traditional FFN layer can be represented as:
$o=W_{0}x+\frac{\alpha}{r}\sum_{i=1}^{N}\omega_{i}\cdot B_{i}A_{i}x$ (5)
@@
where $\omega_{i}$ denotes the weight of $i$-th expert and $\alpha$ is the constant hyper-parameter, approximately equivalent to the learning rate.
@@
Figure 5: The coefficient of variation for the experts of the unconstrained LoRAMoE progressively escalates and sustains at a high value, i.e., approximately three, similar to the phenomenon observed at Shazeer et al. (2016). This indicates that the router assigns large weights to the same few experts.
3.2 Localized Balancing Constraint
The imbalance of the experts’ utilization is a typical problem in MoE (Shazeer et al., 2016; Fedus et al., 2021), which is also observed in our proposed method, as shown in Figure 5. The conventional solution is balancing expert utilization (Shazeer et al., 2016), which involves making the coefficient of variation of the experts’ importance as the loss function. However, this method assumes all the training samples are under the same distribution, which ignores the fact that samples may be from different distributions such as the question-answering task and other downstream tasks, more detailed analysis and conceptual proof in Appendix C.
Considering the mixed characteristics of data distributions are important, during the training phase, we introduce localized balancing constraint, a novelty balancing expert utilization method to make a portion of experts focus more on leveraging world knowledge to solve tasks. As shown in Figure 6, during the fine-tuning phase, we softly constrain experts to concentrate on two aspects, one of which focuses on leveraging world knowledge by learning on its related datasets, while another focuses on other downstream tasks. In addition, all experts within the same aspects are balanced such as balancing expert utilization.
Figure 6: Localized balancing constraint. We softly force experts to focus on two types, one for leveraging world knowledge by learning on its related tasks, and another for concentrating on other downstream tasks. Meanwhile, the experts in solving the same aspect are balancing.